How do I solve the equation of 7x-5+5x+5=-2x+6? It should be converted into a fraction. If possible.
Answer by Alan3354(69443)

How do I solve the equation of 7x-5+5x+5=-2x+6?
-------------
It's an equation. You probably want to solve the equation for x, right?
7x-5+5x+5=-2x+6
Collect terms
12x = -2x+6
Add 2x to both sides
14x = 6
Divide by 14
x = 6/14
x = 3/7
----------
It's important in life, not just in math, to say clearly what you want to do.